# WAPITI: A Watermark for Finetuned Open-Source LLMs

### v1

@@ Abstract
1 Introduction
As large language models (LLMs; Touvron et al., 2023; OpenAI et al., 2024) have been integrated into numerous workflows and play an increasingly significant role in everyday life, controlling these LLMs to prevent potential harm has become even more urgent. Watermarking offers a viable solution by embedding traceable information in model outputs. It enables the identification of LLM-generated content and can be used to trace back to the source model, serving as a methodological foundation for regulatory oversight of language models.
@@ 1 Introduction
Open-source models release their full parameters to users, and users can fully customize the generation process. Therefore, users can simply choose an unwatermarked decoding algorithm to evade watermarking, thereby invalidating existing decoding-based watermarking methods. Gu et al. (2024) proposed a parameter-based method that distills the model using watermarked generations. This process, referred to as watermark distillation, ensures that the watermarks are retained within the model parameters, preventing users from easily removing them.
However, we observe that this method (Gu et al., 2024) would impair the fine-tuned capabilities of models, revealing it is not compatible with fine-tuned models. Additionally, watermark distillation incurs significantly higher computational costs compared to typical fine-tuning. Furthermore, a severe weakness of parameter-based watermarks is their vulnerability to fine-tuning attacks, where malicious users fine-tune the watermarked models with unwatermarked datasets to eliminate their watermarking. As noted by  Gu et al. (2024), as few as 500 steps of fine-tuning attack can remove the watermark from models. (See Table 1 for overall comparison).
To address these limitations, we propose a new train-free watermarking strategy that transfers watermarks from base models to fine-tuned models (WAPITI, WAtermark Parameter InTegratIon) as shown in Figure 1. We discover that watermarking bears a similar effect on the output distribution of both base models and fine-tuned models. The core of our method involves embedding watermarks into models through direct parameter editing, ensuring compatibility with fine-tuned models. Most importantly, WAPITI effectively defends against fine-tuning attacks by binding watermarking with the fine-tuning capabilities of the model.
@@
Figure 1: Previous parameter-based watermarking (left) uses distillation which would impair models’ fine-tuned capabilities. WAPITI (middle) uses watermark-related parameters to transfer watermarking from the base model to fine-tuned models. This method can preserve fine-tuned model capabilities and meanwhile enables them to generate watermarked texts where the green tokens indicate the watermarked tokens (right).
2 Preliminary
2.1 Decoding-based Watermarking
Large Language Models are generally neural networks based on the transformer architecture, denoted as $f_{\bm{\theta}}:\mathcal{V}^{*}\to\Delta(\mathcal{V})$, which maps a given prefix string $\bm{x}\in\mathcal{V}^{*}$ to a probability distribution over the vocabulary $\Delta(\mathcal{V})$ for predicting the next token, denoted as $f_{\bm{\theta}}(\,\cdot\mid\bm{x})$. The generation process involves two main steps: logit generation followed by token sampling (Vaswani et al., 2023).
Method Closed-source Open-sourced Problem
LLMs Base LLMs Fine-tuned LLMs
Decoding-based Easily removable by user
Distillation-based N/A It undermines capabilities
WAPITI (ours) N/A N/A N/A
Table 1: A taxonomy of LLM watermarking. ”N/A” indicates that the method is not designed for the corresponding setting.
Decoding-based watermarks are embedded in either stages of generation with the aim of guiding the output distribution toward a targeted direction, incorporating traceable information for detection. For instance, KGW (Kirchenbauer et al., 2024a) increases the frequency of specific tokens during the generation process, and the detector identifies the origin of a text based on the occurrence rate of these tokens. More specifically, a watermarking algorithm $\mathcal{W}$ employs a watermark key $\phi$ to modify the original next-token distribution $f_{\bm{\theta}}(\,\cdot\mid\bm{x})$ into a watermarked version. The watermark detector $\mathcal{D}$, using the same watermark key $\phi$, can then retrieve the embedded watermark information. In general, given a text $x$ and a watermark key $\phi$, the detector $\mathcal{D}$ calculates a p-value for the null hypothesis that the text $x$ is unrelated to $\mathcal{W}$ and $\phi$. A text is classified as model-generated if its p-value falls below a predefined threshold.
@@ 2.1 Decoding-based Watermarking
Logit-based: KGW is a watermarking strategy applied directly to output logits of the model (Algorithm 2 in Kirchenbauer et al. (2024a)). During the next token generation, the vocabulary is pseudorandomly split into green and red lists based on the previous $k$ tokens. When $k=0$ (Zhao et al., 2023), the green and red lists are fixed, and when $k\geq 1$, the lists are determined by the previous context. The green list contains $\gamma\in(0,1)$ proportion of the entire vocabulary, and an additional watermark shift $\delta$ is added to the logits of the tokens in the green list. This increases the probability of the green tokens being selected in the final generation. During detection, the p-value is calculated by checking whether the proportion of green list tokens exceeds the predefined $\gamma$.
@@
Sampling-based: AAR is the Gumbel softmax scheme from Aaronson (2023), which is a special sampling strategy. When generating $x_{i}$, it hashes the previous $k$ tokens using the key $\phi$ to generate a pseudorandom score sequence $\bm{r}_{i}$ for the entire vocabulary $\mathcal{V}$ where $\bm{r}_{i}\in\mathbb{R}^{|\mathcal{V}|}$ whose entries are uniformly distributed in $[0,1]$. Given the probability distribution $\bm{p}_{i}\in\Delta(\mathcal{V})$ of the next token $x_{i}$, AAR uses Gumbel-Max sampling strategy: $x_{i}=\operatorname*{arg\,max}_{j\in\mid\mathcal{V}\mid}(\log p_{i,j}-\log(-\log r_{i,j}))$ (Cane & Luce, 1960), which introduces some randomness into the sampling stage by adding Gumbel noise $\bm{r}_{i}$. This sampling strategy would result in watermarked texts having comparative higher score sums. During detection, a larger score sum corresponds to a lower p-value against the null hypothesis.
2.2 Weight-based Watermarking
@@
Since the weights of open-source models are fully released, users can modify the decoding method or apply any post-processing to the logits, making decoding-based watermarks easy to remove. The most feasible approach [^2] To the best of our knowledge, this is the only approach for watermarking open-source LLMs that cannot be easily removed by users. for watermarking is to embed the watermark into the model parameters, enabling LLMs to generate watermarked text under natural sampling distribution. Current research (Gu et al., 2024) has shown that LLMs can learn watermarks via distillation and generate detectable watermarked texts. By using decoding-based watermark strategies to generate watermarked texts as distillation data, Gu et al. (2024) has verified the learnability of multiple watermarks on Llama-2-7B and Pythia-1.4B models. However, we found that this parameter-based method is specifically designed for base LLMs. In the fine-tuning setting, it significantly impairs the fine-tuned capabilities, as we will demonstrate in § 3.1.
3 Method
3.1 Motivating Study
@@
Limitation of current weight-based watermarking.  The current weight-based method enables the base model to generate watermarked texts via distillation. In this paper, we explore whether the distillation-based approach is compatible with fine-tuned models. Specifically, we ask: can watermark distillation retain the fine-tuned capabilities of model while embedding the watermark into the fine-tuned model? To address this question, we conduct a preliminary experiment.
In this experiment, we use math fine-tuned Llama-2-7B and decoding-based watermarking strategies to generate watermarked math-related data, which will be used in watermark distillation. Detailed experiment setup can be found in Appendix A.
@@
Figure 3 compares the watermark detectability (measured by p-value) and fine-tuning utility between the original fine-tuned model and the watermark-distilled models. The utility of the model on GSM8K drops sharply to nearly zero, and the output text shows poor detectability, with a p-value close to the baseline of 0.5.
@@
To better understand this phenomenon, we further analyze the distillation data generated by the math-fine-tuned model and identify two main reasons: (i) The quality of the watermarked math data is inferior to that of the original fine-tuning dataset. Table 3 presents several samples from the original benchmark dataset alongside the answers generated by the math model. Although the generated answers might still be correct, they often contain flawed procedures or random repetitive sequences. Such data can confuse the model and result in a performance decline. (ii) The quantity of watermarked math data is insufficient for the student model to learn the watermark effectively. As noted in  Gu et al. (2024), approximately 1.3 million samples are required for a distilled model to internalize the watermark. With only 7.3k samples in the GSM8K training split and further filtering due to the 40% accuracy of model, our final dataset was just 0.6% of the required size. The insufficient distillation data makes the generated outputs lack detectability.
@@
Universal distribution shift from watermarking.  The primary issue with the current weight-based method is the distillation phase, which underscores the need for a train-free approach to watermark fine-tuned models. To this end, we aim to investigate whether there are any similarities between the base models and fine-tuned models when watermarked.
Figure 2: We distill upon a math fine-tuned Llama-2-7B model. Results show that watermark distillation not only fails to embed the watermark but also significantly impairs the math capability of model.
Figure 3: JS divergence between watermarked and unwatermarked $n$-grams across different sifting thresholds. Results show that watermarking has partially aligned the output distributions of the base model and fine-tuned model.
@@
The results, shown in Figure 3, indicate that the JS divergence is consistently smaller for watermarked $n$-gram compared to unwatermarked $n$-gram, which suggests that the distribution of watermarked $n$-gram is more similar between base models and fine-tuned models. This indicates that watermarking distorts the output of both the base and fine-tuned models in similar ways by increasing the frequency of watermarked $n$-grams in the final generation.
3.2 Watermark Parameter
In this section, we focus on deriving the watermarked parameters of fine-tuned models. As mentioned in §2.1, watermarks only perturb the next-token generation $x_{t}$ according to previous $k$ tokens $x_{t-k},\cdots,x_{t-1}$ and watermark key $\phi$, so that watermark perturbation in next-token probability $f_{\bm{\theta}}(\bm{x})$ [^3] For brevity, we identify the next-token probability predictor $f_{\bm{\theta}}(\,\cdot\mid\bm{x}):\mathcal{V}\to\mathbb{R}$ as a vector $f_{\bm{\theta}}(\bm{x})\in\Delta(\mathcal{V})$. remains the same across different models, where $\bm{x}$ is the input prompt. We denote the watermark perturbation as $\delta\cdot g(\bm{x})$, where $\delta$ represents the intensity of the shift, analogous to the watermark shift $\delta$ in KGW and $g(\bm{x})$ is analogous to the mask of green list in KGW watermarking that indicates which part of vocabulary will be applied watermark shift. According to experiments in C, we observe that model parameters can learn watermarking. Let $\bm{\theta}_{\text{Base}},\bm{\theta}_{\text{Base}}^{\dagger}$ represent parameters of the base model and the watermark-distilled base model respectively. So we have:
$f_{\bm{\theta_{\text{Base}}}^{\dagger}}(\bm{x})=f_{\bm{\theta_{\text{Base}}}}(\bm{x})+\delta_{\text{Base}}\cdot g(\bm{x}).$ (1)
Similarly, we use $\bm{\theta}_{\text{FT}}$ and $\bm{\theta}_{\text{FT}}^{\dagger}$ to represent the parameters of the fine-tuned (FT) models, as well as its watermark-distilled counterpart respectively. Our ultimate goal is, given an unwatermarked $\bm{\theta}$, to find the parameter $\bm{\theta}_{\text{FT}}^{\dagger}$ such that:
$f_{\bm{\theta}_{\text{FT}}^{\dagger}}(\bm{x})=f_{\bm{\theta}_{\text{FT}}}(\bm{x})+\delta_{\text{FT}}\cdot g(\bm{x}),$ (2)
@@ 3.2 Watermark Parameter
Let $\Delta\bm{\theta_{\text{Base}}}:=\bm{\theta_{\text{Base}}}^{\dagger}-\bm{\theta_{\text{Base}}}$ and $\Delta\bm{\theta_{\text{FT}}}:=\bm{\theta_{\text{FT}}}^{\dagger}-\bm{\theta_{\text{FT}}}$ denote the parameter differences introduced by watermark distillation for the base and fine-tuned models, respectively. We can eliminate $g(\bm{x})$ by substituting $\bm{\theta}_{\text{Base}},\bm{\theta}_{\text{Base}}^{\dagger}$ into Eq. (1) and rearranging it as a Taylor expansion. $\delta_{\text{Base}}$ denotes the watermark shift of base model:
$g(\bm{x})=\frac{1}{\delta_{\text{Base}}}\left(f_{\bm{\theta}_{\text{Base}}^{\dagger}}(\bm{x})-f_{\bm{\theta}_{\text{Base}}}(\bm{x})\right)=\frac{1}{\delta_{\text{Base}}}\langle\nabla_{\bm{\theta}}f_{\bm{\theta_{\text{Base}}}}(\bm{x}),\Delta\bm{\theta}_{\text{Base}}\rangle+O(\|\Delta\bm{\theta}_{\text{Base}}\|^{2}).$ (3)
Furthermore, we observe in Appendix E.3, that the parameter difference between the fine-tuned model and the base model, $\bm{\theta_{\text{FT}}}-\bm{\theta_{\text{Base}}}$, is approximately orthogonal to the parameter difference caused by watermarking, $\bm{\theta_{\text{Base}}}^{\dagger}-\bm{\theta_{\text{Base}}}$:
$\langle\bm{\theta}_{\text{FT}}-\bm{\theta}_{\text{Base}},\bm{\theta}_{\text{Base}}^{\dagger}-\bm{\theta}_{\text{Base}}\rangle\approx 0.$ (4)
Let $\otimes$ denote the tensor product between differentiation operators, and let ×$\times_{1},\times_{2}$ denote the mode-1 and mode-2 tensor–matrix product, respectively. Let $\bm{H}_{\text{Base}}(\bm{x}):=\nabla_{\bm{\theta}}\otimes\nabla_{\bm{\theta}}f_{\bm{\theta}_{\text{Base}}}(\bm{x})$ be the Hessian. As shown in prior studies, every channel of $\bm{H}_{\text{Base}}(\bm{x})$ is approximately the identity matrix $\bm{I}$ (Jiao et al., 2024; Yang et al., 2024). Combining it with our observation in Eq. (4), we hypothesize that:
$\bm{H}_{\text{Base}}(\bm{x})\times_{1}(\bm{\theta}_{\text{FT}}-\bm{\theta}_{\text{Base}})\times_{2}(\bm{\theta}_{\text{Base}}^{\dagger}-\bm{\theta}_{\text{Base}})\approx 0.$ (5)
The first-order Taylor expansion of $\nabla_{\bm{\theta}}f_{\bm{\theta}_{\text{FT}}}(\bm{x})$ around $\bm{\theta}=\bm{\theta}_{\text{Base}}$ is:
$\nabla_{\bm{\theta}}f_{\bm{\theta}_{\text{FT}}}(\bm{x})=\nabla_{\bm{\theta}}f_{\bm{\theta}_{\text{Base}}}(\bm{x})+\bm{H}_{\text{Base}}(\bm{x})\times_{1}(\bm{\theta}_{\text{FT}}-\bm{\theta}_{\text{Base}})+O(\|\bm{\theta}_{\text{FT}}-\bm{\theta}_{\text{Base}}\|^{2}),$ (6)
$\bm{H}_{\text{Base}}(\bm{x})\times_{1}(\bm{\theta}_{\text{FT}}-\bm{\theta}_{\text{Base}})\approx\nabla_{\bm{\theta}}f_{\bm{\theta}_{\text{FT}}}(\bm{x})-\nabla_{\bm{\theta}}f_{\bm{\theta}_{\text{Base}}}(\bm{x}).$ (7)
@@
$(\nabla_{\bm{\theta}}f_{\bm{\theta}_{\text{FT}}}(\bm{x})-\nabla_{\bm{\theta}}f_{\bm{\theta}_{\text{Base}}}(\bm{x}))\Delta\bm{\theta}_{\text{Base}}\approx 0.$ (8)

### v2

@@ Abstract.
Watermark, Large Language Model, Model Intervention
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Digital rights management
Figure 1. Previous parameter-based watermarking (left) uses distillation which would impair models’ fine-tuned capabilities. WAPITI (middle) uses watermark-related parameters to transfer watermarking from the base model to fine-tuned models. This method can preserve fine-tuned model capabilities and meanwhile enables them to generate watermarked texts where the green tokens indicate the watermarked tokens (right).
1. Introduction
As large language models (LLMs) (Touvron et al., 2023; OpenAI et al., 2024) have been integrated into numerous workflows and play an increasingly significant role in everyday life, controlling these LLMs to prevent potential harm has become even more urgent. Watermarking offers a viable solution by embedding traceable information in model outputs. It enables the identification of LLM-generated content and can be used to trace back to the source model, serving as a methodological foundation for regulatory oversight of language models.
@@ 1. Introduction
Open-source models release their full parameters to users, and users can fully customize the generation process. Therefore, users can simply choose an unwatermarked decoding algorithm to evade watermarking, thereby invalidating existing decoding-based watermarking methods. (Gu et al., 2024) proposed a parameter-based method that distills the model using watermarked generations. This process, referred to as watermark distillation, ensures that the watermarks are retained within the model parameters, preventing users from easily removing them.
However, we observe that this method (Gu et al., 2024) would impair the fine-tuned capabilities of models, revealing it is not compatible with fine-tuned models. Additionally, watermark distillation incurs significantly higher computational costs compared to typical fine-tuning. Furthermore, a severe weakness of parameter-based watermarks is their vulnerability to fine-tuning attacks, where malicious users fine-tune the watermarked models with unwatermarked datasets to eliminate their watermarking. As noted by  (Gu et al., 2024), as few as 500 steps of fine-tuning attack can remove the watermark from models. (See Table 1 for overall comparison).
To address these limitations, we propose a new train-free222”Train-free” means that applying WAPITI to fine-tuned models does not involve any additional model training after acquiring the watermark parameter. watermarking strategy that transfers watermarks from base models to fine-tuned models (WAPITI, WAtermark Parameter InTegratIon) as shown in Figure 1. We discover that watermarking bears a similar effect on the output distribution of both base models and fine-tuned models. The core of our method involves embedding watermarks into models through direct parameter editing, ensuring compatibility with fine-tuned models. Most importantly, WAPITI effectively defends against fine-tuning attacks by binding watermarking with the fine-tuning capabilities of the model.
@@
2. Preliminary
2.1. Decoding-based Watermarking
Large Language Models are generally neural networks based on the transformer architecture, denoted as $f_{\boldsymbol{\theta}}:\mathcal{V}^{*}\to\Delta(\mathcal{V})$, which maps a given prefix string $\boldsymbol{x}\in\mathcal{V}^{*}$ to a probability distribution over the vocabulary $\Delta(\mathcal{V})$ for predicting the next token, denoted as $f_{\boldsymbol{\theta}}(\,\cdot\mid\boldsymbol{x})$. The generation process involves two main steps: logit generation followed by token sampling (Vaswani et al., 2023).
Method Closed-source Open-sourced Open-sourced Application
LLMs Base LLMs Fine-tuned LLMs Efficiency Vulnerability Attack Type
Decoding-based $\mathcal{O}\left(1\right)$ Transparent to users Directly remove watermark modules
Distillation-based N/A $\mathcal{O}\left(\mathcal{C}_{FT}\right)$ Watermark decay with finetuning Finetuning attack
WAPITI N/A $\mathcal{O}\left(\mathcal{C}_{FT}/N\right)$ None None
Table 1. A taxonomy of LLM watermarking. ”N/A” indicates that the method is not designed for the corresponding setting. And $\mathcal{C}_{FT}$ indicates the computation cost of watermark distillation. $N$ indicates the number of models of the same type in that WAPITI only requires one watermark distillation to watermark all models of the same architecture.
Decoding-based watermarks are embedded in either stages of generation with the aim of guiding the output distribution toward a targeted direction, incorporating traceable information for detection. For instance, KGW (Kirchenbauer et al., 2024a) increases the frequency of specific tokens during the generation process, and the detector identifies the origin of a text based on the occurrence rate of these tokens. More specifically, a watermarking algorithm $\mathcal{W}$ employs a watermark key $\phi$ to modify the original next-token distribution $f_{\boldsymbol{\theta}}(\,\cdot\mid\boldsymbol{x})$ into a watermarked version. The watermark detector $\mathcal{D}$, using the same watermark key $\phi$, can then retrieve the embedded watermark information. In general, given a text $x$ and a watermark key $\phi$, the detector $\mathcal{D}$ calculates a p-value for the null hypothesis that the text $x$ is unrelated to $\mathcal{W}$ and $\phi$. A text is classified as model-generated if its p-value falls below a predefined threshold.
@@ 2.1. Decoding-based Watermarking
Logit-based: KGW is a watermarking strategy applied directly to output logits of the model (Algorithm 2 in (Kirchenbauer et al., 2024a)). During the next token generation, the vocabulary is pseudorandomly split into green and red lists based on the previous $k$ tokens. When $k=0$ (Zhao et al., 2023), the green and red lists are fixed, and when $k\geq 1$, the lists are determined by the previous context. The green list contains $\gamma\in(0,1)$ proportion of the entire vocabulary, and an additional watermark shift $\delta$ is added to the logits of the tokens in the green list. This increases the probability of the green tokens being selected in the final generation. During detection, the p-value is calculated by checking whether the proportion of green list tokens exceeds the predefined $\gamma$.
@@
Sampling-based: AAR is the Gumbel softmax scheme from (Aaronson, 2023), which is a special sampling strategy. When generating $x_{i}$, it hashes the previous $k$ tokens using the key $\phi$ to generate a pseudorandom score sequence $\boldsymbol{r}_{i}$ for the entire vocabulary $\mathcal{V}$ where $\boldsymbol{r}_{i}\in\mathbb{R}^{|\mathcal{V}|}$ whose entries are uniformly distributed in $[0,1]$. Given the probability distribution $\boldsymbol{p}_{i}\in\Delta(\mathcal{V})$ of the next token $x_{i}$, AAR uses Gumbel-Max sampling strategy: $x_{i}=\operatorname*{\arg\max}_{j\in\mid\mathcal{V}\mid}(\log p_{i,j}-\log(-\log r_{i,j}))$ (Cane and Luce, 1960), which introduces some randomness into the sampling stage by adding Gumbel noise $\boldsymbol{r}_{i}$. This sampling strategy would result in watermarked texts having comparative higher score sums. During detection, a larger score sum corresponds to a lower p-value against the null hypothesis.
2.2. Weight-based Watermarking
@@
Since the weights of open-source models are fully released, users can modify the decoding method or apply any post-processing to the logits, making decoding-based watermarks easy to remove. The most feasible approach [^3] To the best of our knowledge, this is the only approach for watermarking open-source LLMs that cannot be directly removed by users. for watermarking is to embed the watermark into the model parameters, enabling LLMs to generate watermarked text under natural sampling distribution. Current research (Gu et al., 2024) has shown that LLMs can learn watermarks via distillation and generate detectable watermarked texts. By using decoding-based watermark strategies to generate watermarked texts as distillation data, (Gu et al., 2024) has verified the learnability of multiple watermarks on Llama-2-7B and Pythia-1.4B models. However, we found that this parameter-based method is specifically designed for base LLMs. In the fine-tuning setting, it significantly impairs the fine-tuned capabilities, as we will demonstrate in § 3.1.
3. Method
3.1. Motivating Study
Limitation of current weight-based watermarking.
@@ 3.1. Motivating Study
The current weight-based method enables the base model to generate watermarked texts via distillation. In this paper, we explore whether the distillation-based approach is compatible with fine-tuned models. Specifically, we ask: can watermark distillation retain the fine-tuned capabilities of the model while embedding the watermark into the fine-tuned model? To address this question, we conduct a preliminary experiment.
To obtain a watermarked fine-tuned model using watermark distillation, there are three possible approaches: (i) Distilling a fine-tuned model with watermarked content, (ii) Fine-tuning a distilled model that already contains a watermark, or (iii) Fine-tuning a base model using a watermarked fine-tuning dataset. We use math-fine-tuned Llama-2-7B and decoding-based watermarking strategies to obtain a watermarked math model. To counteract the low-entropy nature of mathematical tasks, we use Chain-of-Thought (CoT) reasoning, allowing the model to derive the final answer step by step. This approach increases the entropy space and enhances the model’s ability to detect previous watermarking. Detailed experimental setups can be found in Appendix A.
@@
Figure 2 (Right) compares the watermark detectability (measured by the negative logarithm of p-value) and fine-tuning utility of the resulting model from all three different approaches. The utility of the models on GSM8K drops sharply to nearly zero, and the output text shows poor detectability, with a p-value close to the baseline of 0.5.
To better understand this phenomenon, we further analyze the three approaches. The first two methods both involve two-phase fine-tuning, which, as studied in previous research, can lead to capability degradation or catastrophic forgetting (Wang et al., 2023).
@@
For the third method, we believe it holds the most potential to enable the base model to learn mathematical capabilities while embedding the watermark content. Therefore, we focus on analyzing the distillation data generated by the math-fine-tuned model and identify two main reasons: (i) The quality of the watermarked math data is inferior to that of the original fine-tuning dataset. Although the generated answers might still be correct, they often contain flawed procedures or random repetitive sequences. Such data can confuse the model and result in a performance decline. (ii) The quantity of watermarked math data is insufficient for the student model to learn the watermark effectively. As noted in  (Gu et al., 2024), approximately 320K samples are required for a distilled model to internalize the watermark. With only 7.3k samples in the GSM8K training split and further filtering due to the 40% accuracy of the model, our final dataset was just 2.2% of the required size. The insufficient distillation data makes the generated outputs lack detectability.
@@
Universal distribution shift from watermarking.
The primary issue with the current weight-based method is the distillation phase, which underscores the need for a train-free approach to watermark fine-tuned models. To this end, we aim to investigate whether there are any similarities between the base models and fine-tuned models when watermarked.
Figure 2. Performance of watermark distillation on fine-tuned models (left) and the distribution alignment of watermarked content compared with unwatermarked content (right) measured with JS divergence.
@@
The results, shown in Figure 2 (Left), indicate that the JS divergence is consistently smaller for watermarked $n$-gram compared to unwatermarked $n$-gram, which suggests that the distribution of watermarked $n$-gram is more similar between base models and fine-tuned models. This indicates that watermarking distorts the output of both the base and fine-tuned models in similar ways by increasing the frequency of watermarked $n$-grams in the final generation.
3.2. Watermark Parameter
In this section, we focus on deriving the watermarked parameters of fine-tuned models. As mentioned in §2.1, watermarks only perturb the next-token generation $x_{t}$ according to previous $k$ tokens $x_{t-k},\cdots,x_{t-1}$ and watermark key $\phi$, so that watermark perturbation in next-token probability $f_{\boldsymbol{\theta}}(\boldsymbol{x})$ [^4] For brevity, we identify the next-token probability predictor $f_{\boldsymbol{\theta}}(\,\cdot\mid\boldsymbol{x}):\mathcal{V}\to\mathbb{R}$ as a vector $f_{\boldsymbol{\theta}}(\boldsymbol{x})\in\Delta(\mathcal{V})$. remains the same across different models, where $\boldsymbol{x}$ is the input prompt. We denote the watermark perturbation as $\delta\cdot g(\boldsymbol{x})$, where $\delta$ represents the intensity of the shift, analogous to the watermark shift $\delta$ in KGW and $g(\boldsymbol{x})$ is analogous to the mask of green list in KGW watermarking that indicates which part of vocabulary will be applied watermark shift. According to experiments in C, we observe that model parameters can learn watermarking. Let $\boldsymbol{\theta}_{\text{Base}},\boldsymbol{\theta}_{\text{Base}}^{\dagger}$ represent parameters of the base model and the watermark-distilled base model respectively. So we have:
(1) $f_{\boldsymbol{\theta_{\text{Base}}}^{\dagger}}(\boldsymbol{x})=f_{\boldsymbol{\theta_{\text{Base}}}}(\boldsymbol{x})+\delta_{\text{Base}}\cdot g(\boldsymbol{x}).$
Similarly, we use $\boldsymbol{\theta}_{\text{FT}}$ and $\boldsymbol{\theta}_{\text{FT}}^{\dagger}$ to represent the parameters of the fine-tuned (FT) models, as well as its watermark-distilled counterpart respectively. Our ultimate goal is, given an unwatermarked $\boldsymbol{\theta}$, to find the parameter $\boldsymbol{\theta}_{\text{FT}}^{\dagger}$ such that:
(2) $f_{\boldsymbol{\theta}_{\text{FT}}^{\dagger}}(\boldsymbol{x})=f_{\boldsymbol{\theta}_{\text{FT}}}(\boldsymbol{x})+\delta_{\text{FT}}\cdot g(\boldsymbol{x}),$
@@ 3.2. Watermark Parameter
Let $\Delta\boldsymbol{\theta_{\text{Base}}}:=\boldsymbol{\theta_{\text{Base}}}^{\dagger}-\boldsymbol{\theta_{\text{Base}}}$ and $\Delta\boldsymbol{\theta_{\text{FT}}}:=\boldsymbol{\theta_{\text{FT}}}^{\dagger}-\boldsymbol{\theta_{\text{FT}}}$ denote the parameter differences introduced by watermark distillation for the base and fine-tuned models, respectively. We can eliminate $g(\boldsymbol{x})$ by substituting $\boldsymbol{\theta}_{\text{Base}},\boldsymbol{\theta}_{\text{Base}}^{\dagger}$ into Eq. (1) and rearranging it as a Taylor expansion. $\delta_{\text{Base}}$ denotes the watermark shift of base model:
(3) $g(\boldsymbol{x})$ $=\frac{1}{\delta_{\text{Base}}}\left(f_{\boldsymbol{\theta}_{\text{Base}}^{\dagger}}(\boldsymbol{x})-f_{\boldsymbol{\theta}_{\text{Base}}}(\boldsymbol{x})\right)$
(4) $=\frac{1}{\delta_{\text{Base}}}\langle\nabla_{\boldsymbol{\theta}}f_{\boldsymbol{\theta_{\text{Base}}}}(\boldsymbol{x}),\Delta\boldsymbol{\theta}_{\text{Base}}\rangle+O(\|\Delta\boldsymbol{\theta}_{\text{Base}}\|^{2}).$
Furthermore, we observe in Appendix 4.7, that the parameter difference between the fine-tuned model and the base model, $\boldsymbol{\theta_{\text{FT}}}-\boldsymbol{\theta_{\text{Base}}}$, is approximately orthogonal to the parameter difference caused by watermarking, $\boldsymbol{\theta_{\text{Base}}}^{\dagger}-\boldsymbol{\theta_{\text{Base}}}$:
(5) $\langle\boldsymbol{\theta}_{\text{FT}}-\boldsymbol{\theta}_{\text{Base}},\boldsymbol{\theta}_{\text{Base}}^{\dagger}-\boldsymbol{\theta}_{\text{Base}}\rangle\approx 0.$
Let $\otimes$ denote the tensor product between differentiation operators, and let ×$\times_{1},\times_{2}$ denote the mode-1 and mode-2 tensor–matrix product, respectively. Let $\boldsymbol{H}_{\text{Base}}(\boldsymbol{x}):=\nabla_{\boldsymbol{\theta}}\otimes\nabla_{\boldsymbol{\theta}}f_{\boldsymbol{\theta}_{\text{Base}}}(\boldsymbol{x})$ be the Hessian. As shown in prior studies, every channel of $\boldsymbol{H}_{\text{Base}}(\boldsymbol{x})$ is approximately the identity matrix $\boldsymbol{I}$ (Jiao et al., 2024; Yang et al., 2024). Combining it with our observation in Eq. (5), we hypothesize that:
(6) $\boldsymbol{H}_{\text{Base}}(\boldsymbol{x})\times_{1}(\boldsymbol{\theta}_{\text{FT}}-\boldsymbol{\theta}_{\text{Base}})\times_{2}(\boldsymbol{\theta}_{\text{Base}}^{\dagger}-\boldsymbol{\theta}_{\text{Base}})\approx 0.$
The first-order Taylor expansion of $\nabla_{\boldsymbol{\theta}}f_{\boldsymbol{\theta}_{\text{FT}}}(\boldsymbol{x})$ around $\boldsymbol{\theta}=\boldsymbol{\theta}_{\text{Base}}$ is:
$\nabla_{\boldsymbol{\theta}}f_{\boldsymbol{\theta}_{\text{FT}}}(\boldsymbol{x})$ $=\nabla_{\boldsymbol{\theta}}f_{\boldsymbol{\theta}_{\text{Base}}}(\boldsymbol{x})+\boldsymbol{H}_{\text{Base}}(\boldsymbol{x})\times_{1}(\boldsymbol{\theta}_{\text{FT}}-\boldsymbol{\theta}_{\text{Base}})$
(7) $\quad+O(\|\boldsymbol{\theta}_{\text{FT}}-\boldsymbol{\theta}_{\text{Base}}\|^{2})$
(8) $\boldsymbol{H}_{\text{Base}}(\boldsymbol{x})\times_{1}(\boldsymbol{\theta}_{\text{FT}}-\boldsymbol{\theta}_{\text{Base}})\approx\nabla_{\boldsymbol{\theta}}f_{\boldsymbol{\theta}_{\text{FT}}}(\boldsymbol{x})-\nabla_{\boldsymbol{\theta}}f_{\boldsymbol{\theta}_{\text{Base}}}(\boldsymbol{x}).$
@@
(9) $(\nabla_{\boldsymbol{\theta}}f_{\boldsymbol{\theta}_{\text{FT}}}(\boldsymbol{x})-\nabla_{\boldsymbol{\theta}}f_{\boldsymbol{\theta}_{\text{Base}}}(\boldsymbol{x}))\Delta\boldsymbol{\theta}_{\text{Base}}\approx 0.$